In a certain​ region, the probability of selecting an adult over 40 years of age with a certain disease is 0.03 If the probability of correctly diagnosing a person with this disease as having the disease is 0.90 and the probability of incorrectly diagnosing a person without the disease as having the disease is 0.01​, what is the probability that an adult over 40 years of age is diagnosed with the​ disease?
Answer #1

Answer:

Given that:

In a certain​ region, the probability of selecting an adult over 40 years of age with a certain disease is 0.03 If the probability of correctly diagnosing a person with this disease as having the disease is 0.90 and the probability of incorrectly diagnosing a person without the disease as having the disease is 0.01​

probability of diagnosed with disease =P(have disease and diagnosed with +do not have disease and diagnosed with )

= 0.03*0.9+(1-0.03)*0.01

= 0.027 + 0.97 * 0.01

= 0.027 + 0.0097

= 0.0367

NOTE:

If you are satisfied with my answer please do upvote and if you have any kind of doubts please post in comment section. i'll surely help you there.
Thank You:)